A company imports a product from China at a cost of $50 per unit. The product is imported with an import tariff of 20% and it is sold for $100 per unit. The total sales revenues for the company were 85.000. If the government increases the import tariff per unit of the product to 30% and the elasticity of demand is 3 find the following:


(i) How many units of this product the company will sell if the price rises the same amount with the import tariff.

(ii) Calculate if the revenues of the company will increase or decrease with the increase of import tariff.


1
Expert's answer
2022-02-05T06:58:49-0500

(i)

"E=\frac{\Delta Q}{\Delta P}"


"\Delta P=\frac{110}{100}=1.1"


"3=\frac{\Delta Q}{1.1}"


"\Delta Q=3.3"

"3.3=\frac{Q1}{850}"


Q1=2805


(ii)

"2805\times110-850\times100=223550"

increase